# Quantitative approaches for multi-scale structural analysis with atomic resolution electron microscopy

### v1

@@ I Introduction
1 Since the development of aberration correction, atomic-resolution imaging with scanning transmission electron microscopy (STEM) has been increasingly depended upon as an invaluable part of the materials characterization tool set. In addition to sub-Angstrom spatial resolution, atomic column positions can be measured with picometer-scale precision, and sensitivity to composition or local lattice strain can be tuned with high- and low-angle collection geometries [1, 2]. More recently, broad in situ capacity, including high resolution imaging at variable temperatures and under applied external fields, makes possible direct correlation of material properties with atomic-scale structural changes, while developing techniques such as phase retrieval with electron ptychography promise to extend these capabilities even further [3, 4, 5, 6].
2 In the STEM, a condensed (typically $<$1 Å) electron beam is scanned across a thin crystalline sample oriented along a high-symmetry axis to resolve columns of atoms in projection along the beam direction. The resulting atomic-resolution images contain rich information about spatial variations in material composition, symmetry, and defects encoded in the image contrast. These images are exceptionally useful in studying structural features such as local fields including polarization and strain [7, 8], broken symmetries including those associated with charge or chemical order [9, 10], as well as the presence of various crystalline phases, domains, defects, and other sources of disorder [11]. A key challenge, however, is disentangling the particular features of interest from the large amount of structural information present in STEM images, as well as in the presence of measurement noise and imaging artifacts.
@@
3 A variety of analysis approaches have been developed to handle the many distinct structural features of interest (e.g., inter-planar spacings, inter-atomic column angles, magnitudes of displacements, etc., as depicted in Figure 1). These approaches range from simple operations, such as extracting a profile of image intensity along a lattice plane, to complex algorithms and associated specialized software packages such as pycroscopy [12, 13], Atomap [14], TopoTEM [15], motif-learn [16, 17], and TEMMETA [18]. While some analyses can be accomplished entirely in real space, other approaches benefit from working in reciprocal space, which can be accessed by taking the Fourier transform of the STEM image. Reciprocal space offers a convenient basis for measuring symmetries and the spacing and orientation of lattice planes, while the separation of features based on spatial frequency can be taken advantage of to map specific periodicities back to real space [19, 20, 21, 9]. For instance, spatial variations in an inter-planar spacing might be measured in real space by calculating the distances between sets of atomic columns, or with Fourier based geometric phase analysis (GPA) [7, 20, 8]. Working in real space and reciprocal space each have advantages and disadvantages in terms of measurement precision and resolution, as well as different levels of susceptibility to noise, computational efficiency, and other trade-offs. Here, we present two approaches that synergize real and reciprocal space analysis taking advantage of the distinct benefits of working in both spaces.
4 These image processing algorithms address two key STEM image analysis problems: (1) measuring translational symmetry-breaking structural distortions, and (2) quantifying variations in lattice parameters and local strains. Our Fourier damping approach for characterizing symmetry breaking can be applied to a wide range of distortions including periodic lattice displacements, octahedral tilting, and chemical or vacancy orders, is robust in the presence of defects, strain and inhomogeneity, and can disentangle coexisting distortions [9]. Meanwhile lattice parameter measurements based on real-space wave fitting to Fourier filtered signals offer interpretable quantification of inter-planar spacings and other parameters on an absolute scale, as well as an understanding of the associated uncertainty in the measurements [22]. While here we focus on high-angle annular dark-field (HAADF) STEM imaging, these approaches are also applicable to other STEM modalities, ptychographic reconstructions, and conventional high-resolution TEM analysis, and can be extended to other high-resolution real space characterization techniques such as scanning probe microscopies. The algorithms and a suite of atomic-resolution image analysis tools are available in the open source kemstem Python package at https://github.com/noahschnitzer/kemstem.
@@ II Mapping symmetry breaking with Fourier damping
5 Material functionality and properties of interest are directly tied to crystal symmetry and symmetry breaking structural distortions [23]. For instance, polar distortions which break inversion symmetry are necessary for ferroelectricity, coupling of charge and orbital order to the lattice break rotation and translational symmetries, and manipulating mirror and inversion symmetries can tune photonic properties [23, 24, 25, 26]. The symmetry group of a distorted structure will be a subgroup of a “parent” undistorted higher symmetry structure, so the natural basis in which to characterize a broken symmetry is with reference to the parent structure. For STEM image analysis, this poses the problem of constructing a high symmetry reference for comparison with the experimental measurement – generally not a trivial task. In some cases, a directly measurable feature of interest such as a change in inter-column distances or angles can be used as an ad hoc surrogate, but in many cases no such proxy is available. We will show that for translational symmetry breaking distortions – those with a non-zero wavevector - damping the associated peaks in reciprocal space is a systematically applicable method for generating a high symmetry reference to enable the mapping of structural distortions in STEM images [9, 27].
@@
6 Breaking translational symmetry forms a superlattice with a supercell larger than the parent unit cell, and a corresponding reciprocal space lattice with smaller reciprocal lattice wavevectors than the parent. Thus, in the Fourier transform of a STEM image, the structural information associated with a translational symmetry breaking distortion will be localized to frequencies at the reciprocal lattice points associated with the superstructure. A portion of these superlattice points will be distinct from the reciprocal lattice points of the parent, forming satellite peaks at linear combinations of the distortion wavevectors around the parent lattice points (e.g. as marked in cyan and pink in Fig. 1) creating a convenient partitioning in reciprocal space between the distorted and parent structures.
@@
7 Taking advantage of this feature separation in reciprocal space, a high-symmetry reference can be generated from an experimental image by damping the amplitude of the superlattice peaks in its Fourier transform to the background level - without altering the phase - and taking the inverse Fourier transform to change back to a real space basis. In the resulting false image, the symmetry breaking distortion will be suppressed, resembling the undistorted parent structure. Because the Fourier transform amplitude is only altered at the superlattice points and the phase is left unchanged, other changes to the experimental STEM image will be minimal. Features such as interfaces, boundaries, defects, strains, and even other symmetry breaking distortions with different wavevectors will be preserved in the resulting reference image.
@@
8 Following the workflow described in Ref. [9], we demonstrate in Figure 2 how this approach enables the measurement of picometer-scale periodic lattice displacements in the manganese oxide SmBaMn2O6 . SmBaMn2O6 crystallizes with a double perovskite structure where samarium and barium alternate on the A-site along [001] [28]. When cooled the material develops symmetry breaking charge order and associated atomic displacements [29]. Figure 2a shows an atomic resolution HAADF-STEM image acquired on the [001] zone axis: the A-site alternation is lost in projection, but the positions of the atomic columns are modulated by displacements (too small to see by eye) associated with charge order, which give rise to superlattice peaks in the image’s Fourier transform (Fig. 2b). The magnitude of the Fourier transform is damped to the background level in circular areas around each superlattice peak while the phase is left unchanged (Fig. 2c). Taking the real part of the inverse Fourier transform results in an image almost identical to the original, but with minute shifts in the positions of the atomic columns towards their undistorted, high-symmetry positions (Fig. 2d). Mapping out the change in the image contrast by taking the difference in the intensity of each pixel between the original image and the high-symmetry reference reveals a periodic pattern of dumbbell-like features centered on the atomic columns (Fig. 2e). These dumbbells arise from shifts of the atomic column positions smaller than the widths of the columns themselves, which result in reduced intensity on one side of the column and increased intensity on the other. These shifts are quantified by fitting 2D Gaussian functions to each atomic column in both the original and reference image to precisely determine their positions. The vectors from the position of each column in the reference to the position in the original image describe the atomic displacements associated with charge order in the material, and can be mapped out to visualize the structural changes in the material attributable to the broken symmetry (Fig. 2f).
@@
9 This approach has been extensively applied to mapping the atomic displacements associated with charge order [9, 30, 19, 31, 32], as well as antipolar displacements [33, 34], and even artifacts from channeling of the electron beam [35], but it can also be readily generalized to characterizing other kinds of symmetry breaking that do not result in a clear displacement of atomic columns. Any periodic symmetry breaking with a non-zero wavevector (i.e. which generates superlattice peaks) that alters the image contrast can be characterized with this technique, meaning it can be applied to chemical and vacancy order, as well as distortions which may not be coherent along the beam direction [36]. Such distortions can be recognized by the absence of the characteristic dumbbell pattern associated with atomic column displacements seen in Figure 2e, and while similar image analysis can be applied it requires different interpretation and visualization than that shown in Figure 2. Figure 3 demonstrates these distinctions with an analysis of the chemical order also present in SmBaMn2O6 . Viewed along the [100] zone axis, alternation of the A-site cation between samarium ($Z=62$) and barium ($Z=56$) creates a layered structure (Fig. 3b), with a small but discernible modulation in the atomic number dependent contrast of HAADF-STEM imaging (Fig. 3a). This ordering can be identified by taking line profiles along the A-site cation columns (Fig. 3c), but visualizing inhomogeneity in the chemical order requires mapping the associated intensity modulations across full STEM images. Extensive analysis of local intensity profiles would be tedious at best, and at worst computationally intractable to do over large areas with sufficient real space sensitivity.
@@
10 The alternation of samarium and barium columns doubles the size of the unit cell along the $c$-axis, generating superlattice peaks in reciprocal space with a wavevector of $\frac{1}{2}c^{*}$ (Fig. 3d). As in the case of the periodic lattice displacements induced by charge order, masking these peaks results in a reference image where the symmetry breaking distortion - in this case, the modulation in $A$-site contrast - is suppressed. Figure 3e shows the difference between the original STEM image and this high symmetry reference, which reveals the layers of alternating contrast in the A-site planes, where blue indicates higher intensity (i.e., Sm) and red a lower intensity (i.e., Ba), but lacks the dumbbell features associated with atomic column displacements seen previously (Fig. 2e). Mapping out the chemical order in this way also reveals an anti-phase boundary in the order, clearly visible extending vertically in the difference map. The presence of this defect can be confirmed by comparing intensity profiles measured on either side of the STEM image which show opposite intensity modulations on the $A$-site columns (Fig. 3c). By mapping distortions continuously in real space with Fourier damping, however, defects can be identified without prior knowledge of their presence and inhomogeneity over large length scales can be systematically measured.
@@
11 For robust, scalable analysis of STEM images, understanding the effects of parameter choices on the image processing is critical. For the Fourier damping approach described here, two key parameters are: (1) the number of superlattice peaks damped, and (2) the size of the region damped around each peak. Varying the number of peaks damped controls what information is mapped back to real space and can affect the accuracy of measured distortions (Suppl. Fig. S2). The presence of inhomogeneity, including defects like the anti-phase boundary shown in Figure 3, will broaden the superlattice peaks associated with symmetry-breaking distortions. The information in the Fourier transform which encodes the morphology of the order will extend away from the reciprocal lattice points into their peak tails, meaning that the size of the damped reciprocal space regions is also an important parameter to consider when conducting this analysis.
@@
12 The quantitative dependence of Fourier damping measurements on these parameters – size and number of damping masks – is assessed on simulated STEM images with known atomic displacements. Varying both the number of peaks ($N$) and the radius of the circular region of reciprocal space about each peak ($\sigma$) which are damped, displacements are first measured on a test image of a structure modulated by uniform sinusoidal atomic displacements similar to those shown in Figure 2. Damping fewer superlattice peaks reduces the measured amplitude of the atomic displacements, as can be seen in Figure 4a. As the number of damped peaks increases, the measured displacement amplitude approaches – but remains systematically smaller than – that of the true displacements in the structure (Fig. 4b). Thus, damping a large set of peaks including those at large angles and higher order satellites - in practice, as many superlattice peaks as can be identified in the Fourier transform - is critical to estimate the amplitude of atomic displacements, and such measurements should be interpreted as a lower bound on the true displacement.
@@
13 On the other hand, for a uniform modulation changing the damping radius has little effect, unless it is made so small that most of the peak magnitude is not masked at all, or so large that other peaks, such as the neighboring Bragg peaks, are also damped. In either of these extremes the resulting displacements are entirely incorrect, but for a large range of intermediate values the displacements are accurately measured (Fig. 4c,d). When the distortion incorporates significant inhomogeneity, however, the damping radius becomes more significant. The top panel of Figure 4e shows a model structure with an atomically sharp interface between an area with periodic lattice displacements and an area without. As in the case of a uniform modulation, damping more or fewer superlattice peaks alters the amplitude of the measured displacements, but has little effect on the sharpness of the interface (Fig. 4e,f). On the other hand, adjusting the damping radius has a strong effect on the sharpness with which the interface is resolved. As can be seen in Figure 4g, when only a very small area around the peak center is damped, the interface is not resolved at all. Rather, distorted and undistorted domains are only hinted by a gradual ramp in the displacement amplitude. As the damping radius is increased, the spatial resolution is enhanced and the measurement quickly approaches the model structure. Over a broad range of radii ($\sigma=0.34q$ to $0.51q$) the measurement is sharp and accurate to the model distortion, indicating that even for inhomogenous distortions a wide range of parameters will provide reliable measurement. Care must nonetheless still be taken that the radius is not made so large that other peaks (or even their tails) are included within it, as this introduces serious artifacts and makes the measurement uninterpretable. While for the simple model presented here this effect is not apparent until an extremely large radius impinges directly on the Bragg peak ($\sigma=1.02$), when applying the technique attention should be paid to what features are damped and various damping radii should be tested for consistency. Full two-dimensional sweeps through both parameters are included in Supplementary Note 3 [37].
@@
14 Line profiles of the displacement magnitude across the interfaces make the trends even more clear - at low damping radii, the displacement profiles resemble a Fourier series with very few terms. As more pixels of the Fourier transform are damped, the series converges towards accurately representing the sharp interface (Fig. 4f,h). This result can be understood by considering the delocalization of information in reciprocal space in the presence of inhomogeneity: a real-valued plane wave involves only a single spatial frequency or one Fourier coefficient, and in reciprocal space it is represented by a conjugate pair of delta functions. When the wavelength and amplitude of the wave are allowed to spatially vary, however, the wave will extend over a range of frequencies in reciprocal space and require representation with multiple Fourier coefficients. As the sharpness of these features increases, more coefficients are needed for an accurate representation. Consequently, the damping radius is inversely related to coarsening in real space, so large damping radii must be used for accurately measuring spatial variations in inhogmogenous distortions. This highlights a constraint on the technique: large-wavelength or low-frequency distortions will have small wavevectors and little separation between superlattice and Bragg peaks in the Fourier transform, limiting the achievable spatial resolution for such measurements [9].
@@
15 The Fourier damping approach presented here enables robust measurement of symmetry breaking distortions in atomic-resolution images through the generation of self-consistent high-symmetry reference images. By taking advantage of the localization of these distortions in reciprocal space, a wide variety of phenomena including periodic lattice displacements, chemical order, vacancy order, and more can be analyzed, and the measurements directly represent physically meaningful changes in symmetry. Coexisting distortions associated with distinct wavevectors can be decoupled, as demonstrated for instance in Ref. [9]. Clear heuristics for parameter selection are established and benchmarking on model datasets with known distortions demonstrates the high accuracy and precision achievable with this technique.

### v2

@@ I Introduction
Since the development of aberration correction, atomic-resolution imaging with scanning transmission electron microscopy (STEM) has been increasingly depended upon as an invaluable part of the materials characterization tool set. In addition to sub-Angstrom spatial resolution, atomic column positions can be measured with picometer-scale precision, and sensitivity to composition or local lattice strain can be tuned with high- and low-angle collection geometries [1, 2]. More recently, broad in situ capacity, including high resolution imaging at variable temperatures and under applied external fields, makes possible direct correlation of material properties with atomic-scale structural changes, while developing techniques such as phase retrieval with electron ptychography promise to extend these capabilities even further [3, 4, 5, 6].
In the STEM, a condensed (typically <<1 Å) electron beam is scanned across a thin crystalline sample oriented along a high-symmetry axis to resolve columns of atoms in projection along the beam direction. The resulting atomic-resolution images contain rich information about spatial variations in material composition, symmetry, and defects encoded in the image contrast. These images are exceptionally useful in studying structural features such as local fields including polarization and strain [7, 8], broken symmetries including those associated with charge or chemical order [9, 10], as well as the presence of various crystalline phases, domains, defects, and other sources of disorder [11]. A key challenge, however, is disentangling the particular features of interest from the large amount of structural information present in STEM images, as well as in the presence of measurement noise and imaging artifacts.
@@
A variety of analysis approaches have been developed to handle the many distinct structural features of interest (e.g., inter-planar spacings, inter-atomic column angles, magnitudes of displacements, etc., as depicted in Figure 1). These approaches range from simple operations, such as extracting a profile of image intensity along a lattice plane, to complex algorithms and associated specialized software packages such as pycroscopy [12, 13], Atomap [14], TopoTEM [15], motif-learn [16, 17], and TEMMETA [18]. While some analyses can be accomplished entirely in real space, other approaches benefit from working in reciprocal space, which can be accessed by taking the Fourier transform of the STEM image. Reciprocal space offers a convenient basis for measuring symmetries and the spacing and orientation of lattice planes, while the separation of features based on spatial frequency can be taken advantage of to map specific periodicities back to real space [19, 20, 21, 9]. For instance, spatial variations in an inter-planar spacing might be measured in real space by calculating the distances between sets of atomic columns, or with Fourier based geometric phase analysis (GPA) [7, 20, 8]. Working in real space and reciprocal space each have advantages and disadvantages in terms of measurement precision and resolution, as well as different levels of susceptibility to noise, computational efficiency, and other trade-offs. Here, we present two approaches that synergize real and reciprocal space analysis taking advantage of the distinct benefits of working in both spaces.
These image processing algorithms address two key STEM image analysis problems: (1) measuring translational symmetry-breaking structural distortions, and (2) quantifying variations in lattice parameters and local strains. Our Fourier damping approach for characterizing symmetry breaking can be applied to a wide range of distortions including periodic lattice displacements, octahedral tilting, and chemical or vacancy orders, is robust in the presence of defects, strain and inhomogeneity, and can disentangle coexisting distortions [9]. Meanwhile lattice parameter measurements based on real-space wave fitting to Fourier filtered signals offer interpretable quantification of inter-planar spacings and other parameters on an absolute scale, as well as an understanding of the associated uncertainty in the measurements [22]. While here we focus on high-angle annular dark-field (HAADF) STEM imaging, these approaches are also applicable to other STEM modalities, ptychographic reconstructions, and conventional high-resolution TEM analysis, and can be extended to other high-resolution real space characterization techniques such as scanning probe microscopies. The algorithms and a suite of atomic-resolution image analysis tools are available in the open source kemstem Python package at https://github.com/noahschnitzer/kemstem.
@@ II Mapping symmetry breaking with Fourier damping
Material functionality and properties of interest are directly tied to crystal symmetry and symmetry breaking structural distortions [23]. For instance, polar distortions which break inversion symmetry are necessary for ferroelectricity, coupling of charge and orbital order to the lattice break rotation and translational symmetries, and manipulating mirror and inversion symmetries can tune photonic properties [23, 24, 25, 26]. The symmetry group of a distorted structure will be a subgroup of a “parent” undistorted higher symmetry structure, so the natural basis in which to characterize a broken symmetry is with reference to the parent structure. For STEM image analysis, this poses the problem of constructing a high symmetry reference for comparison with the experimental measurement – generally not a trivial task. In some cases, a directly measurable feature of interest such as a change in inter-column distances or angles can be used as an ad hoc surrogate, but in many cases no such proxy is available. We will show that for translational symmetry breaking distortions – those with a non-zero wavevector - damping the associated peaks in reciprocal space is a systematically applicable method for generating a high symmetry reference to enable the mapping of structural distortions in STEM images [9, 27].
@@
Breaking translational symmetry forms a superlattice with a supercell larger than the parent unit cell, and a corresponding reciprocal space lattice with smaller reciprocal lattice wavevectors than the parent. Thus, in the Fourier transform of a STEM image, the structural information associated with a translational symmetry breaking distortion will be localized to frequencies at the reciprocal lattice points associated with the superstructure. A portion of these superlattice points will be distinct from the reciprocal lattice points of the parent, forming satellite peaks at linear combinations of the distortion wavevectors around the parent lattice points (e.g. as marked in cyan and pink in Fig. 1) creating a convenient partitioning in reciprocal space between the distorted and parent structures.
@@
Taking advantage of this feature separation in reciprocal space, a high-symmetry reference can be generated from an experimental image by damping the amplitude of the superlattice peaks in its Fourier transform to the background level - without altering the phase - and taking the inverse Fourier transform to change back to a real space basis. In the resulting false image, the symmetry breaking distortion will be suppressed, resembling the undistorted parent structure. This can be understood in terms of the Fourier shift theorem: damping the peak amplitudes in Fourier space re-weights the image’s phase information, resulting in changes in image contrast and shifts in column positions to effectively ‘reverse’ the distortion and reconstruct the higher symmetry parent structure. Because the Fourier transform amplitude is only altered at the superlattice points and the phase is left unchanged, other changes to the experimental STEM image will be minimal. Features such as interfaces, boundaries, defects, strains, and even other symmetry breaking distortions with different wavevectors will be preserved in the resulting reference image.
@@
Following the workflow described in Ref. [9], we demonstrate in Figure 2 how this approach enables the measurement of picometer-scale periodic lattice displacements in the manganese oxide SmBaMn2O6 . SmBaMn2O6 crystallizes with a double perovskite structure where samarium and barium alternate on the A-site along [001] [28]. When cooled the material develops symmetry breaking charge order and associated atomic displacements [29]. Figure 2a shows an atomic resolution HAADF-STEM image acquired on the [001] zone axis: the A-site alternation is lost in projection, but the positions of the atomic columns are modulated by displacements (too small to see by eye) associated with charge order, which give rise to superlattice peaks in the image’s Fourier transform (Fig. 2b). The magnitude of the Fourier transform is damped to the background level in circular areas around each superlattice peak while the phase is left unchanged (Fig. 2c). Taking the real part of the inverse Fourier transform results in an image almost identical to the original, but with minute shifts in the positions of the atomic columns towards their undistorted, high-symmetry positions (Fig. 2d). Mapping out the change in the image contrast by taking the difference in the intensity of each pixel between the original image and the high-symmetry reference reveals a periodic pattern of dumbbell-like features centered on the atomic columns (Fig. 2e). These dumbbells arise from shifts of the atomic column positions smaller than the widths of the columns themselves, which result in reduced intensity on one side of the column and increased intensity on the other. These shifts are quantified by fitting 2D Gaussian functions to each atomic column in both the original and reference image to precisely determine their positions. The vectors from the position of each column in the reference to the corresponding position in the original image describe the atomic displacements associated with charge order in the material and can be mapped to visualize the structural changes in the material attributable to the broken symmetry (Fig. 2f). Construction of the reference image leverages the clean separation of lattice and superlattice signals in Fourier space. Here that construction is performed by damping the superlattice peaks; similar analysis can also be performed via filtering in Fourier space to select only the superlattice peaks, rather than damping them. This will result in a similar signal to Figure 2e, which can be subtracted from the original image to arrive at a high symmetry reference image as discussed in Supplementary Note IV [30].
@@
The Fourier damping approach has been extensively applied to mapping the atomic displacements associated with charge order [9, 31, 19, 32, 33], as well as antipolar displacements [34, 35], and even artifacts from channeling of the electron beam [36], but it can also be readily generalized to characterizing other kinds of symmetry breaking that do not result in a clear displacement of atomic columns. Any periodic symmetry breaking with a non-zero wavevector (i.e. which generates superlattice peaks) that alters the image contrast can be characterized with this technique, meaning it can be applied to chemical and vacancy order, as well as distortions which may not be coherent along the beam direction [37]. Such distortions can be recognized by the absence of the characteristic dumbbell pattern associated with atomic column displacements seen in Figure 2e, and while similar image analysis can be applied it requires different interpretation and visualization than that shown in Figure 2. Figure 3 demonstrates these distinctions with an analysis of the chemical order also present in SmBaMn2O6 . Viewed along the [100] zone axis, alternation of the A-site cation between samarium ($Z=62$) and barium ($Z=56$) creates a layered structure (Fig. 3b), with a small but discernible modulation in the atomic number dependent contrast of HAADF-STEM imaging (Fig. 3a). This ordering can be identified by taking line profiles along the A-site cation columns (Fig. 3c), but visualizing inhomogeneity in the chemical order requires mapping the associated intensity modulations across full STEM images. Extensive analysis of local intensity profiles would be tedious at best, and at worst computationally intractable to do over large areas with sufficient real space sensitivity.
@@
The alternation of samarium and barium columns doubles the size of the unit cell along the $c$-axis, generating superlattice peaks in reciprocal space with a wavevector of $\frac{1}{2}c^{*}$ (Fig. 3d). As in the case of the periodic lattice displacements induced by charge order, masking these peaks results in a reference image where the symmetry breaking distortion - in this case, the modulation in $A$-site contrast - is suppressed. Figure 3e shows the difference between the original STEM image and this high symmetry reference, which reveals the layers of alternating contrast in the A-site planes, where blue indicates higher intensity (i.e., Sm) and red a lower intensity (i.e., Ba), but lacks the dumbbell features associated with atomic column displacements seen previously (Fig. 2e). Mapping out the chemical order in this way also reveals an anti-phase boundary in the order, clearly visible extending vertically in the difference map. The presence of this defect can be confirmed by comparing intensity profiles measured on either side of the STEM image which show opposite intensity modulations on the $A$-site columns (Fig. 3c) or comparing fit intensities of each atomic column in the image (Suppl. Fig. 3) [30]. By mapping distortions continuously in real space with Fourier damping, however, defects can be identified without prior knowledge of their presence and inhomogeneity over large length scales can be systematically measured.
@@
For reliable analysis, understanding the effects of parameter choices on the image processing is critical. Here, a set of superlattice peaks are identified and the amplitude of circular regions of the Fourier transform around each peak are set uniformly to the local background level. Because the Fourier transform background level varies at different wavevectors, we first fit each peak with a 2D Gaussian function, and set the circular damping area to the fitted background. This leaves two key remaining parameters for the analysis: (1) the number of superlattice peaks damped, and (2) the size of the region damped around each peak. Varying the number of peaks damped controls what information is mapped back to real space and can affect the accuracy of measured distortions (Suppl. Fig. 2) [30]. The presence of inhomogeneity, including defects like the anti-phase boundary shown in Figure 3, will broaden the superlattice peaks associated with symmetry-breaking distortions. The information in the Fourier transform which encodes the morphology of the order will extend away from the reciprocal lattice points into their peak tails, meaning that the size of the damped reciprocal space regions is also an important parameter to consider when conducting this analysis.
@@
The quantitative dependence of Fourier damping measurements on these parameters – size and number of damping masks – is assessed on simulated STEM images with known atomic displacements. Varying both the number of peaks ($N$) and the radius of the circular region of reciprocal space about each peak ($\sigma$) which are damped, displacements are first measured on a test image of a structure modulated by uniform sinusoidal atomic displacements similar to those shown in Figure 2. Damping fewer superlattice peaks reduces the measured amplitude of the atomic displacements, as can be seen in Figure 4a. As the number of damped peaks increases, the measured displacement amplitude approaches – but remains systematically smaller than – that of the true displacements in the structure (Fig. 4b). Thus, damping a large set of peaks including those at large angles and higher order satellites - in practice, as many superlattice peaks as can be identified in the Fourier transform - is critical to estimate the amplitude of atomic displacements, and such measurements should be interpreted as a lower bound on the true displacement.
@@
On the other hand, for a uniform modulation changing the damping radius has little effect, unless it is made so small that most of the peak magnitude is not masked at all, or so large that other peaks, such as the neighboring Bragg peaks, are also damped. In either of these extremes the resulting displacements are entirely incorrect, but for a large range of intermediate values the displacements are accurately measured (Fig. 4c,d). When the distortion incorporates significant inhomogeneity, however, the damping radius becomes more significant. The top panel of Figure 4e shows a model structure with an atomically sharp interface between an area with periodic lattice displacements and an area without. As in the case of a uniform modulation, damping more or fewer superlattice peaks alters the amplitude of the measured displacements, but has little effect on the sharpness of the interface (Fig. 4e,f). On the other hand, adjusting the damping radius has a strong effect on the sharpness with which the interface is resolved. As can be seen in Figure 4g, when only a very small area around the peak center is damped, the interface is not resolved at all. Rather, distorted and undistorted domains are only hinted by a gradual ramp in the displacement amplitude. As the damping radius is increased, the spatial resolution is enhanced and the measurement quickly approaches the model structure. Over a broad range of radii ($\sigma=0.34q$ to $0.51q$) the measurement is sharp and accurate to the model distortion, indicating that even for inhomogeneous distortions a wide range of parameters will provide reliable measurement. Care must nonetheless still be taken that the radius is not made so large that other peaks (or even their tails) are included within it, as this introduces serious artifacts and makes the measurement uninterpretable. While for the simple model presented here this effect is not apparent until an extremely large radius impinges directly on the Bragg peak ($\sigma=1.02$), when applying the technique attention should be paid to what features are damped and various damping radii should be tested for consistency. Similar analysis of parameter choices on experimental datasets is shown in Supplemental Figure 8 [30]. Full two-dimensional sweeps through both parameters are included in Supplementary Note III [30].
@@
Line profiles of the displacement magnitude across the interfaces make the trends even more clear - at low damping radii, the displacement profiles resemble a Fourier series with very few terms. As more pixels of the Fourier transform are damped, the series converges towards accurately representing the sharp interface (Fig. 4f,h). This result can be understood by considering the delocalization of information in reciprocal space in the presence of inhomogeneity: a real-valued plane wave involves only a single spatial frequency or one Fourier coefficient, and in reciprocal space it is represented by a conjugate pair of delta functions. When the wavelength and amplitude of the wave are allowed to spatially vary, however, the wave will extend over a range of frequencies in reciprocal space and require representation with multiple Fourier coefficients. As the sharpness of these features increases, more coefficients are needed for an accurate representation. Consequently, the damping radius is inversely related to coarsening in real space, so large damping radii must be used for accurately measuring spatial variations in inhomogeneous distortions, just as larger Fourier filters result in a smaller coarsening length (Supplementary Note I) [30]. This highlights a constraint on the technique: large-wavelength or low-frequency distortions will have small wavevectors and little separation between superlattice and Bragg peaks in the Fourier transform, limiting the achievable spatial resolution for such measurements [9].
@@
The Fourier damping approach presented here enables robust measurement of symmetry breaking distortions in atomic-resolution images through the generation of self-consistent high-symmetry reference images. By taking advantage of the localization of these distortions in reciprocal space, a wide variety of phenomena including periodic lattice displacements, chemical order, vacancy order, and more can be analyzed, and the measurements directly represent physically meaningful changes in symmetry. Coexisting distortions associated with distinct wavevectors can be decoupled, as demonstrated for instance in Ref. [9]. Clear heuristics for parameter selection are established and benchmarking on model datasets with known distortions demonstrates the high accuracy and precision achievable with this technique.